# Boosting Active Defense Persistence: A Two-Stage Defense Framework Combining Interruption and Poisoning Against Deepfake

### v2

@@ III-C Perturbation Fusion Optimization
The interruption perturbation and the subsequent poisoning perturbation are then synergistically combined into a unified perturbation. Directly combining the two defensive goals into a single weighted-sum loss function is often suboptimal, as competing gradients can lead to functional conflicts. To avoid this, our intensity separation mechanism decouples the two tasks spatially by utilizing the low-intensity regions of the primary interruption perturbation. This strategy allows both the interruption and poisoning functions to operate effectively without mutual interference, ensuring a more effective and efficient unified perturbation. The process of this mechanism is illustrated in Fig. 4.
@@
The mechanism employs a region-based fusion strategy. To avoid the challenges of a direct multi-objective optimization via a single loss function, this mechanism employs a region-based fusion strategy. It works by first identifying the regions within the initial interruption perturbation $W_{0}$, where the signal magnitude is low. The poisoning mask $M_{d}$, is then generated to isolate these low-intensity areas, designating them as the channels for the poisoning component. Furthermore, an attenuation coefficient $\beta$ scales the poisoning perturbation’s strength, ensuring it is applied precisely where it will not interfere with the high-intensity components of the interruption. we compute the magnitude map:
$M_{d}=e^{-\beta\cdot|W_{0}|}\cdot I({|W_{0}|<\tau}).$ (15)
The exponential term is used to precisely control the allocation of the poisoning perturbation. It ensures that the poisoning is applied most strongly in regions, creating a smooth and effective fusion.
The poisoning perturbation $\delta$ is then optimized within these masked regions $M_{d}$. The final unified perturbation $\delta_{\text{final}}$ is created by adding the poisoning perturbations to the initial interruption perturbation $W_{0}$ and clipping the result to maintain the correct range. This process is iterated to achieve an optimal balance.
$\delta_{t+1}=\text{Clip}(\delta_{t}-\eta\cdot\nabla_{\delta}L_{\text{total}},-\epsilon,\epsilon)\cdot M_{d},$ (16)
@@
where $\tau$ is a predefined threshold and $\eta$ is the step size. This fusion strategy ultimately produces a single perturbation with dual defensive capabilities, the effectiveness of which is quantitatively validated in the following section.
@@
1:Input: Image $x$, feature extractors $E$, detectors $D$, budget $\epsilon$, threshold $\tau$, iterations $T_{int},T_{poi}$.
@@
3:// Stage 1: Generate Interruption Perturbation
4: Initialize interruption perturbation $W_{0}$ randomly;
@@
6:  Update $W_{t}$ using $\mathcal{L}_{\mathrm{feature}}$ (Eq. 9) and Projected Gradient Descent;
7:end for
8: Set $W_{0}\leftarrow W_{T_{int}}$;
9:// Stage 2: Generate and Fuse Poisoning Perturbation
10: Initialize poisoning perturbation $\delta_{poi,0}\leftarrow\mathbf{0}$;
11: Create poisoning mask $M_{d}\leftarrow e^{-\beta\cdot|W_{0}|}\cdot\mathbf{I}(|W_{0}|<\tau)$;
@@
13:  Update $\delta_{poi,t}$ using $\mathcal{L}_{\mathrm{total}}$ (Eq. 14) and the mask $M_{d}$;
14:  Set $\delta_{poison}\leftarrow\delta_{poi,T_{poi}}$;
15:  Combine perturbations: $\delta_{final}\leftarrow W_{0}+\delta_{poison}$;
@@
17:return  $\text{Clip}(\delta_{final},-\epsilon,\epsilon)$;
@@
CelebA StarGAN 0.21 0.26 0.63  100.00% 100.00% 100.00% 305.05 362.54 380.21 12.85 11.45 7.56 0.48 0.26 0.24
AGGAN 0.23 0.19 0.29 99.93% 100.00% 100.00% 163.00 224.97 202.19 15.93 16.12 13.97 0.62 0.68 0.57
AttGAN 0.13 0.15 0.12 87.01% 97.90% 92.55% 180.95 183.22 181.81 18.59 17.36 18.68 0.69 0.64 0.68
HiSD 0.21 0.09 0.17 99.72% 92.55% 98.25% 188.85 183.46 186.81 16.09 19.11 17.66 0.75 0.82 0.79
Avg 0.20 0.17 0.30 96.67% 97.61% 97.70% 209.46 238.55 238.76 15.87 16.01 14.47 0.64 0.60 0.57
LFW StarGAN 0.21 0.26 0.63  100.00% 100.00% 100.00% 298.86 398.86 402.82 12.92 15.86 7.99 0.47 0.27 0.23
AGGAN 0.24 0.20 0.28  100.00% 100.00% 100.00% 193.57 269.06 272.18 16.10 15.72 14.08 0.61 0.66 0.56
AttGAN 0.11 0.16 0.14 95.24% 98.05% 95.50% 207.76 248.99 230.12 18.51 16.70 17.48 0.67 0.60 0.63
HiSD 0.12 0.07 0.08 97.44% 88.04% 91.55% 221.82 163.23 224.17 16.22 19.61 17.63 0.75 0.82 0.79
Avg 0.17 0.17 0.28 98.17% 96.52% 96.75% 230.50 270.04 282.32 15.84 16.97 14.30 0.63 0.59 0.55
FF++O StarGAN 0.21 0.24 0.59 100.00% 100.00% 100.00% 377.31 390.27 424.34 13.58 8.49 13.01 0.55 0.37 0.24
AGGAN 0.24 0.22 0.32 100.00% 100.00% 100.00% 276.09 370.10 304.76 15.14 13.25 15.55 0.70 0.77 0.62
AttGAN 0.14 0.21 0.15 94.80% 99.51% 94.90% 216.97 282.66 240.03 15.71 16.81 15.66 0.70 0.60 0.66
HiSD 0.11 0.09 0.08 86.15% 85.35% 87.75% 213.57 178.71 224.06 17.99 17.10 19.39 0.83 0.88 0.87
Avg 0.18 0.19 0.29 95.24% 96.22% 95.66% 270.99 305.44 298.30 15.61 13.91 16.02 0.70 0.66 0.60
@@
Method Type CelebA LFW FF++O
@@
F1-score $\downarrow$ Original 0.99 0.99 1.00 0.99 0.98 1.00 0.99 0.99 1.00
Random 0.99 0.87 0.99 0.99 0.85 0.99 0.99 0.94 0.99
SSOD 0.81 0.90 0.99 0.81 0.96 0.97 0.90 0.92 0.98
@@ IV-A Experimental Setup
Datasets: We conduct the experiments on the CelebA dataset [31] for training. The dataset comprises over 200,000 face images, from which we use the aligned and cropped version. To ensure a fair comparison with the SOTA baseline FOUND [20] while efficiently learning to disrupt the common and redundant features in facial images, we use a training set of 128 images. For evaluation, we then apply this generated perturbation to all images in the official test split. To further evaluate the generalization capability of TSDF, additional datasets are introduced, including the Labeled Faces in the Wild (LFW) dataset [32] and facial images extracted from the FaceForensics++ Original video dataset (FF++O) [33].
Deepfake Model: We evaluate the defense against four deepfake models, HiSD [34], AttentionGAN (AGGAN) [35], StarGAN [36], and AttGAN [37]. For StarGAN and AGGAN, we select 5 attributes: black hair, blonde hair, brown hair, gender, and age. In AttGAN’s training, the editing tasks are expanded to 14 attributes. Given HiSD’s advanced facial editing capabilities, the experiments incorporated modifications of the glasses attribute. To ensure experimental consistency and fairness, all deepfake models are trained on the CelebA dataset.
@@
Baseline: To rigorously evaluate our framework, our baselines were selected to represent the most foundational, representative and directly competing state-of-the-art (SOTA) methods. For interruption, we compare with the pioneering universal defense method CMUA-Watermark [15] and select FOUND [20] as a critical baseline. The latter’s principle of attacking an ensemble of feature extractors is similar to our own interruption module, making it the most relevant SOTA benchmark to validate our feature enhancement part. For the poisoning task, we benchmark against Face Poison [22], the most direct contemporary competitor that also targets face detection, and the foundational SSOD [23] to validate our effectiveness against detector-level attack strategies. Given the limited availability of public code, these poisoning baselines were rigorously re-implemented from their original papers to ensure a fair comparison.
Implementation Details: The hyperparameter configuration is aligned with existing baseline methods, while jointly optimizing the interruption and poisoning effects of the perturbations. The perturbation range for both attacks is set to (0.05, 0.05) to balance visual fidelity with attack efficacy. In the training stage, 50 and 20 iterations are performed for interruption and poisoning, as we observe in preliminary experiments that performance is saturated beyond these points. For the optimizer parameters in Equations 9 and 10, $\lambda$ and $\gamma$ are set to 0.1 and 0.001, respectively. The intensity separation threshold $\tau$ was set to 0.3, a value determined to provide the optimal balance between defensive functions, as validated in the ablation study.
@@ IV-B Evaluation Metrics
Interruption Metrics: We use L2mask loss and SRmask [20] to measure the success of the interruption. A higher L2mask loss indicates a more successful disruption. The Fréchet Inception Distance (FID) is used to quantify the perceptual difference between the original images and the distorted outputs, with higher values indicating a stronger interruption effect. To assess the disruptive effect, we also employ the Peak Signal-to-Noise Ratio (PSNR) and the Structural Similarity Index (SSIM). Both metrics are used to measure the discrepancy between the original images and the distorted output, where lower scores indicate a better interruption.
@@
Image Quality Metrics: We use PSNR and SSIM to measure the visual quality of the protected images compared to the originals. Higher PSNR and SSIM values indicate that the perturbation is imperceptible. SSIM is used in two distinct contexts: for evaluating interruption, a lower score between the forged and original images indicates a more effective defense. In contrast, to assessing image quality, a higher score indicates better visual imperceptibility.
@@ IV-C Compare the SOTA Methods
This section validates the effectiveness of the proposed TSDF framework through a series of direct comparisons with state-of-the-art (SOTA) methods.
@@
We evaluate the interruption capability of TSDF against SOTA methods, CMUA-Watermark [15] and FOUND [20], with detailed results presented in Table I. The results show that TSDF consistently delivers a more powerful and destructive effect on the final forged images across multiple datasets and models. A key advantage of the method is its superior ability to degrade the structural integrity of the forged output. This is most obvious in SSIM scores. On the CelebA dataset, TSDF achieves an average SSIM of 0.57, which is significantly lower than the scores of CMUA and FOUND, indicating superior disruption. This trend is particularly strong against models like StarGAN, where TSDF achieves a remarkable SSIM of just 0.24, far exceeding the disruptive effect of baseline. Furthermore, this disruptive force is reflected in other key metrics. On the LFW dataset, TSDF achieves the highest average L2mask of 0.28 and the highest average FID score of 282.32, indicating the most powerful disruption among the three methods. The minor difference in some of TSDF’s initial interruption metrics is an expected outcome of our dual-function design. Unlike single-function baselines that dedicate their entire budget to interruption, our framework must balance both interruption and poisoning capabilities within the same perturbation. Our ablation study quantitatively demonstrates that this balance is a deliberate design choice, made to achieve long-term persistence.
@@
Model Metric L2mask↑ SRmask↑ FID↑ PSNR↓ SSIM↓
@@
We then assess the effectiveness of TSDF’s poisoning component against face detectors, comparing it to random noise and reproduced versions of SSOD and FacePoison. The results in Table II are critical as they validate the core defense strategy. Instead of merely trying to survive an attacker’s adaptation, our framework fundamentally disrupts the retraining process. By significantly affecting the automated face detectors that an attacker relies on, our method makes large-scale and automated collection of training data impractical. This represents a fundamental shift in defensive strategy. The objective is no longer to create a perturbation that passively withstands retraining, but to design one that actively destroys the retraining process. This capability is crucial for achieving more persistent defense. For example, on the CelebA dataset, TSDF drastically reduces the F1-score for the RetinaFace detector to 0.78. This represents a far more successful attack compared to FacePoison and SSOD, which only achieve scores of 0.96 and 0.90 respectively. The advantage is even more pronounced against the S3FD detector, where our method achieves an F1-score of 0.72, while the other methods barely impact the detector’s performance. This superior performance is consistent across the LFW and FF++O datasets, where TSDF consistently achieves the lowest F1-score. These results validate that our framework’s poisoning component poses a significant threat to the face detection process. This capability is crucial for our two-stage defense, as it is the mechanism designed to prevent attackers from bypassing the interruption defense.
@@
Poisoned 61.30 0.85
Figure 5: Visual comparison of the interruption effects. This figure compares the outputs of deepfake models when protected by CMUA, FOUND, and TSDF.
@@
TABLE VI: Comparison of SSIM and PSNR indicators on different datasets

### v3

@@ III-C Perturbation Fusion Optimization
The interruption perturbation and the subsequent poisoning perturbation are then synergistically combined into a unified perturbation. Directly combining the two defensive goals into a single weighted-sum loss function is often suboptimal, as competing gradients can lead to functional conflicts [32]. To avoid this, our intensity separation mechanism decouples the two tasks spatially by utilizing the low-intensity regions of the primary interruption perturbation. This strategy allows both the interruption and poisoning functions to operate effectively without mutual interference, ensuring a more effective and efficient unified perturbation. The process of this mechanism is illustrated in Fig. 4.
@@
To avoid the challenges of a direct multi-objective optimization via a single loss function, this mechanism employs a region-based fusion strategy. It works by first identifying the regions within the initial interruption perturbation $W_{0}$, where the signal magnitude is low. The poisoning mask $M_{d}$, is then generated to isolate these low-intensity areas, designating them as the channels for the poisoning component. Furthermore, an attenuation coefficient $\beta$ scales the poisoning perturbation’s strength, ensuring it is applied precisely where it will not interfere with the high-intensity components of the interruption. we compute the magnitude map:
$M_{d}=e^{-\beta\cdot|W_{0}|}\cdot I({|W_{0}|<\tau}),$ (15)
where $\tau$ is a predefined threshold and the exponential term $e^{-\beta\cdot|W_{0}|}$ functions as a principled soft-assignment strategy for adversarial budget allocation. Theoretically, the magnitude of $|W_{0}|$ represents the local sensitivity required to maintain the interruption effect. By employing exponential decay, the framework ensures that poisoning energy is adaptively concentrated in regions with minimal interruption signals, effectively navigating the trade-off between the two defensive functions. Compared to a rigid binary mask, this continuous scaling offers a more fine-grained optimization landscape within the region. By assigning weights based on magnitude rather than a uniform binary value, it provides informative gradients that guide the optimizer to prioritize less sensitive regions. This mechanism mitigates the optimization rigidity often associated with discrete masking, reducing potential functional conflicts during the iterative fusion process.
The poisoning perturbation $\delta$ is then optimized within these masked regions $M_{d}$. Since this stage requires precise convergence within the residuals, we employ a constrained iterative optimization strategy. The update rule is formulated as:
$\delta_{t+1}=\text{Clip}(\delta_{t}-\eta\cdot\nabla_{\delta}L_{\text{total}},-\epsilon,\epsilon)\odot M_{d},$ (16)
@@
where $\eta$ is the step size and $\odot$ denotes element-wise multiplication. This fusion strategy ultimately produces a single perturbation with dual defensive capabilities. The complete workflow of our proposed framework is summarized in Algorithm 1.
@@
1:Input: image $x$, budget $\epsilon$, threshold $\tau$, initial perturbation W, poisoning perturbation $\delta$, iterations $T_{int},T_{poi}$.
@@
3:// Stage 1: Generate Interruption Perturbation.
4: Initialize interruption perturbation $W_{0}$ randomly.
@@
6:  Update W using $\mathcal{L}_{\mathrm{feature}}$ (Eq. 9) via MI-FGSM to maximize.
7:  Set $W_{0}\leftarrow W_{T_{int}}$.
8:end for
9:// Stage 2: Generate and Fuse Poisoning Perturbation.
10: Initialize poisoning perturbation $\delta_{0}\leftarrow\mathbf{0}$.
11: Create poisoning mask $M_{d}\leftarrow e^{-\beta\cdot|W_{0}|}\cdot\mathbf{I}(|W_{0}|<\tau)$.
@@
13:  Update $\delta$ using $\mathcal{L}_{\mathrm{total}}$ (Eq. 14) and the mask $M_{d}$.
14:  Set $\delta_{poison}\leftarrow\delta_{T_{poi}}$.
15:  Combine perturbations: $\delta_{final}\leftarrow W_{0}+\delta_{poison}$.
@@
17:return $\text{Clip}(\delta_{final},-\epsilon,\epsilon)$.
@@
CelebA StarGAN 0.21 0.26 0.37 100.00% 100.00% 100.00% 305.05 362.54 378.89 12.85 11.45 9.73 0.48 0.26 0.17
AGGAN 0.23 0.19 0.18 99.93% 100.00% 100.00% 163.00 224.97 221.36 15.93 16.12 17.86 0.62 0.68 0.70
AttGAN 0.13 0.15 0.17 87.01% 97.90% 98.25% 180.95 183.22 192.31 18.59 17.36 16.76 0.69 0.64 0.62
HiSD 0.21 0.09 0.17 99.72% 92.55% 99.15% 188.85 183.46 170.99 16.09 19.11 16.14 0.75 0.82 0.76
Avg 0.20 0.17 0.22 96.67% 97.61% 99.35% 209.46 238.55 240.88 15.87 16.01 15.12 0.64 0.60 0.56
LFW StarGAN 0.21 0.26 0.36 100.00% 100.00% 100.00% 298.86 398.86 372.80 12.92 15.86 10.32 0.47 0.27 0.20
AGGAN 0.24 0.20 0.19 100.00% 100.00% 100.00% 193.57 269.06 228.19 16.10 15.72 17.97 0.61 0.66 0.69
AttGAN 0.11 0.16 0.16 95.24% 98.05% 99.01% 207.76 248.99 253.76 18.51 16.70 16.17 0.67 0.60 0.59
HiSD 0.12 0.07 0.12 97.44% 88.04% 95.42% 221.82 163.23 188.98 16.22 19.61 16.58 0.75 0.82 0.76
Avg 0.17 0.17 0.21 98.17% 96.52% 98.61% 230.50 270.04 260.93 15.84 16.97 15.26 0.63 0.59 0.56
FF++O StarGAN 0.21 0.24 0.35 100.00% 100.00% 100.00% 377.31 390.27 400.15 13.58 8.49 10.74 0.55 0.37 0.23
AGGAN 0.24 0.22 0.19 100.00% 100.00% 100.00% 276.09 370.10 299.18 15.14 13.25 17.66 0.70 0.77 0.74
AttGAN 0.14 0.21 0.16 94.80% 99.51% 99.02% 216.97 282.66 260.94 15.71 16.81 15.47 0.70 0.60 0.57
HiSD 0.11 0.09 0.11 86.15% 85.35% 94.95% 213.57 178.71 176.98 17.99 17.10 16.67 0.83 0.88 0.81
Avg 0.18 0.19 0.20 95.24% 96.22% 98.49% 270.99 305.44 284.31 15.61 13.91 15.14 0.70 0.66 0.59
@@
Metric Method CelebA LFW FF++O
@@
[…]
Random 0.99 0.85 0.99 0.99 0.85 0.99 0.99 0.84 0.99
SSOD 0.81 0.90 0.99 0.81 0.96 0.97 0.80 0.92 0.98
@@ IV-A Experimental Setup
Datasets: We conduct the experiments on the CelebA dataset [33] for training. The dataset comprises over 200,000 face images, from which we use the aligned and cropped version. To ensure a fair comparison with the SOTA baseline FOUND [27], we adopt a training set of 128 images. As shown in FOUND, the perturbation performance saturates at this sample size meaning 128 images are sufficient to capture the common features for constructing a universal attack. For evaluation, we then apply this generated perturbation to all images in the official test split. To further evaluate the generalization capability of TSDF, additional datasets are introduced, including the Labeled Faces in the Wild (LFW) dataset [34] and facial images extracted from the FaceForensics++ Original video dataset (FF++O) [35].
Deepfake Model: We evaluate the defense against four deepfake models, HiSD [36], AttentionGAN (AGGAN) [37], StarGAN [38], and AttGAN [39]. For StarGAN and AGGAN, we select 5 attributes: black hair, blonde hair, brown hair, gender, and age. In AttGAN’s training, the editing tasks are expanded to 14 attributes. Given HiSD’s advanced facial editing capabilities, the experiments incorporate modifications of the glasses attribute. To ensure experimental consistency and fairness, all deepfake models are trained on the CelebA dataset.
@@
Baseline: To rigorously evaluate our framework, our baselines are selected to represent the most foundational, representative and directly competing SOTA methods. For interruption, we compare with the pioneering universal defense method CMUA-Watermark [19] and select FOUND [27] as a critical baseline. The latter’s principle of attacking an ensemble of feature extractors is similar to our own interruption module, making it the most relevant SOTA benchmark to validate our feature enhancement part. For the poisoning task, we benchmark against Face Poison [22], the most direct contemporary competitor that also targets face detection, and the foundational SSOD [26] which provides a representative baseline for disrupting the fundamental classification and localization mechanisms within the face detection pipeline to validate our effectiveness against detector-level attack strategies. Given the limited availability of public code, these poisoning baselines are rigorously re-implemented from their original papers to ensure a fair comparison.
Implementation Details: The hyperparameter configuration is aligned with existing baseline methods while jointly optimizing the interruption and poisoning effects of the perturbations. The perturbation range for both attacks is set to (0.05, 0.05) to balance visual fidelity with attack efficacy. In the training stage, 50 and 30 iterations are performed for interruption and poisoning, as we observe in preliminary experiments that performance is saturated beyond these points. For the optimizer parameters in Equation 10 and 15, $\lambda$ and $\beta$ are set to 0.001 and 0.4, respectively. The intensity separation threshold $\tau$ is set to 0.04, a value determined to provide the optimal balance between defensive functions, as validated in the ablation study. For the interruption stage utilizing MI-FGSM, the momentum decay factor $\mu$ is set to 0.8.
@@ IV-B Evaluation Metrics
Interruption Metrics: We use L2mask value and SRmask [27] to measure the success of the interruption. A higher L2mask value indicates a more successful disruption. The Fréchet Inception Distance (FID) is used to quantify the perceptual difference between the original images and the distorted outputs, with higher values indicating a stronger interruption effect. To assess the disruptive effect, we also employ the Peak Signal-to-Noise Ratio (PSNR) and the Structural Similarity Index Measure(SSIM). Both metrics are used to measure the discrepancy between the original images and the distorted output, where lower scores indicate a better interruption.
@@
Image Quality Metrics: We use PSNR and SSIM to measure the visual quality of the protected images compared to the originals. Higher PSNR and SSIM values indicate that the perturbation is imperceptible. PSNR and SSIM are used in two distinct contexts: for evaluating interruption, a lower score between the forged and original images indicates a more effective defense. In contrast, to assessing image quality, a higher score indicates better visual imperceptibility.
@@ IV-C Compare the SOTA Methods
This section validates the effectiveness of the proposed TSDF method through a series of direct comparisons with SOTA methods.
@@
We evaluate the interruption capability of TSDF against SOTA methods, CMUA-Watermark [19] and FOUND [27], with detailed results presented in Table I. The results show that TSDF consistently delivers a more powerful and destructive effect on the final forged images across multiple datasets and models. A key advantage of the method is its superior ability to degrade the structural integrity of the forged output. This is most obvious in SSIM scores. On the CelebA dataset, TSDF achieves an average SSIM of 0.56, which is significantly lower than the scores of CMUA and FOUND, indicating superior disruption. This trend is particularly strong against models like StarGAN, where TSDF achieves a remarkable SSIM of just 0.17, far exceeding the disruptive effect of baseline. Furthermore, this disruptive force is reflected in other key metrics. On the LFW dataset, TSDF achieves the highest average L2mask of 0.21 and the highest average SRmask of 98.61%, indicating the most powerful disruption among the three methods. The minor difference in some of TSDF’s initial interruption metrics is an expected outcome of our dual-function design. Unlike single-function baselines that dedicate their entire budget to interruption, our framework must balance both interruption and poisoning capabilities within the same perturbation. Our ablation study quantitatively demonstrates that this balance is a deliberate design choice, made to achieve long-term persistence.
@@
Method Deepfake Model L2mask↑ SRmask↑ FID↑ PSNR↓ SSIM↓
@@
We then assess the effectiveness of TSDF’s poisoning component against face detectors, comparing it to random noise and reproduced versions of SSOD and FacePoison. The results in Table II are critical as they validate the core defense strategy. Instead of merely trying to survive an attacker’s adaptation, our framework fundamentally disrupts the retraining process. By significantly affecting the automated face detectors that an attacker relies on, our method makes large-scale and automated collection of training data impractical. This represents a fundamental shift in defensive strategy. The objective is no longer to create a perturbation that passively withstands retraining, but to design one that actively destroys the retraining model. This capability is crucial for achieving more persistent defense. For example, on the CelebA dataset, TSDF drastically reduces the F1-score for the RetinaFace detector to 0.77. This represents a far more successful attack compared to FacePoison and SSOD, which only achieve scores of 0.96 and 0.90 respectively. The advantage is even more pronounced against the S3FD detector, where our method achieves F1-score of 0.23, while the other methods barely impact the detector’s performance. This superior performance is consistent across the LFW and FF++O datasets, where TSDF consistently achieves the lowest F1-score. These results validate that our framework’s poisoning component poses a significant threat to the face detection process. This capability is crucial for our two-stage defense, as it is the mechanism designed to prevent attackers from bypassing the interruption defense.
@@
Poisoned 121.06 0.68
Figure 5: Visual comparison of the interruption effect. This figure compares the outputs of deepfake models when protected by CMUA, FOUND, and TSDF.
@@
TABLE VI: Comparison of SSIM and PSNR indicators on different datasets.